K*BUG Booth Guide Documentation

Release

K*BUG Booth Guide

Feb 26, 2017

CONTENTS

1 Daemon-Mushi (でもんむし) 3 2 Interest 5 3 Meeting 7 4 Event activities 9 5 AsiaBSDCon 2017 11 6 Kansai * BSD Users Group 2017/2/11 meeting 13

6.1 Presentations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 7 Open Source Conference (OSC) 2017 Osaka 17

7.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 8 KOF 2016 19

8.1 BSD BOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 9 OSC2016 Kyoto 21

9.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 10 OSC2015 Kansai@Kyoto 27

10.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 11 KOF 2014 31

11.1 BSD BOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 12 OSC2014 Kansai@Kyoto 33

12.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 13 OSC2013 Kansai@Kyoto 35

13.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 14 Event @ Nara National College of Technology(NNCT) 13rd 39 15 KOF 2012 43

15.1 BSD BoF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 16 OSC2012 Kansai@Kyoto 49

16.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

\overline{i} 17 Event @ NNCT 12nd 53 18 KOF 2011 55

18.1 BSD BoF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 19 OSC2011 Kansai@Kyoto 57

19.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 20 OSC2010 Kansai@Kyoto 61

20.1 Booth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 21 OSC2010 Kansai@Kobe 63

21.1 JNUG Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 22 KOF 2009 65 23 OSC2009 Kansai@Kyoto 67

23.1 Booth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 24 Nico-Tech: Kyoto Meeting (NK:M) 69 25 KOF 2008 71 26 OSC2008 Kansai@Kyoto 75 27 KOF 2007 81 28 Event @ NNCT 11st 83 29 KOF 2006 87 30 Event @ NNCT 3rd 89

30.1 Presentations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 31 Event @ NNCT 1st 91

\overline{ii}

K*BUG Booth Guide Documentation, Release

Kansai * BSD Users Group (K*BUG) is regional BSD Users Group at Kansai area in Japan. Kansai area has many cities including Osaka, Kyoto, Kobe, Nara and so on. Our web page is http://www.kbug.gr.jp/ .

Fig. 1: Kansai area (c) OpenStreetMap contributors CHAPTER

ONE

DAEMON-MUSHI (でもんむし)

Deamon-Mushi (でもんむし) is a mascot character for K*BUG. Inkscape SVG source file is at http://sacraya.610t.org/kbug/logo/ .

CHAPTER

TWO

INTEREST

• BSD: FreeBSD (PC-BSD), NetBSD, OpenBSD, DragonlyBSD, (macOS?), (iOS??)...

• arch: i386, amd64, arm (Raspberry Pi), macppc, landisk, zaurus, wzero3, hpcmips, netwalker, ipaq, fonera, vax, ...

• kernel hack: kernel module, BLE, serial driver, ...

• New technology: DTrace, ZFS, ...

• Software package: FreeBSD ports, pkgsrc, ...

• Security: pkg audit, maintainance my server,

• Education: Scratch, Squeak, ...

• Languages: Prolog, Lisp, awk, Squeak, Scratch, ...

• Hardware: UPS, HDD, Server, Bluetooth, GPIO, ...

• Physical Computing: Gainer (gainerm-lib), Arduino, ...

• Electronic work: hacked USB audio servo controller, electric circuit with magazine, ...

• misc: OpenCV, ContaoCMS, ...

• Woodworking: K*BUG sign, daemon wood carving, ...

• Drinking party after every meeting CHAPTER

THREE

MEETING

We hold meeting every 2 monthes. Meeting schedule is http://www.kbug.gr.jp/workshops.html . We use these meeting rooms:

• AXE Inc. meeting room at Kyoto: http://www.axe-inc.co.jp/english/index.html

• Grand front Osaka: https://www.grandfront-osaka.jp/multilingulal/ CHAPTER

FOUR

EVENT ACTIVITIES

We open booth at these events: • Open Source Conference : http://ospn.jp/

• Kansai Open Forum : https://k-of.jp/

• Event@NNCT : http://qml.610t.org/NNCT/event.html CHAPTER

FIVE

ASIABSDCON 2017

• Official page: https://2017.asiabsdcon.org/

• 2017/3/9(Wed)-12(Sun) at Tokyo University of Science

• Facebook: https://www.facebook.com/events/1718621151740857/ • togetter: https://togetter.com/li/1084357 CHAPTER

SIX

KANSAI * BSD USERS GROUP 2017/2/11 MEETING

• Official page : http://www.kbug.gr.jp/events/2017-1st.html

• 2017/2/11(Sat) at Axe inc (Kyoto)

• Facebook : https://www.facebook.com/events/920179428083721/ • togetter : https://togetter.com/li/1078967

K*BUG Booth Guide Documentation, Release • RetroBSD: 2.11BSD on MIPS(PIC32MX). The board name is Pinguino-RetroBSD.

• LiteBSD: 4.4BSDLite on MIPS(PIC32MZ). The board name is Chipkit WiFire. Serial Console is Pocket C.H.I.P.

• Raspberry Pi Zero

6.1 Presentations

• IOzone graph with Excel

• K*BUG @ OSC2015 : http://qml.610t.org/FreeBSD/OSC2015Kyoto.html

• Raspberry Pi 3 on FreeBSD

• Scratch on BSD : http://qml.610t.org/FreeBSD/ScratchBSD.html

• iOS with wired LAN CHAPTER

SEVEN

OPEN SOURCE CONFERENCE (OSC) 2017 OSAKA

• Official page : https://www.ospn.jp/osc2017-osaka/

• 2017/1/27(Fri), 28(Sat) at Osaka Sansokan

• K*BUG flyer (not published) : http://sacraya.610t.org/kbug/flyer/flyer-OSC2017Osaka.pdf

• togetter by JNUG : https://togetter.com/li/1075115

• netbsd-advocacy: “NetBSD machines at Open Source Conference 2017 Osaka” http://mail-index.netbsd.org/netbsd-advocacy/2017/01/28/msg000727.html

7.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2017osaka.pdf

• PC6001  emulator  PC6001VX  on  NetBSD  +  pkgsrc  by  @tsutsui  :

https://speakerdeck.com/tsutsui/osc2017-osaka-pc-6001-emulator-on-netbsd-and-pkgsrc CHAPTER

EIGHT

KOF 2016

• Official page : https://k-of.jp/2016/

• 2016/11/11(Fri), 12(Sat) at ATC (Osaka)

• togetter by JNUG : https://togetter.com/li/1047263

• netbsd-advocacy: “NetBSD machines at KANSAI OPEN FORUM 2016” http://mail-index.netbsd.org/netbsd-advocacy/2016/11/16/msg000722.html

8.1 BSD BOF

• NetBSD travel guide : http://www.soum.co.jp/~jun/KOF2016.pdf

• NetBSD/dreamcast IDE HDD again (more update from OSC2016Tokyo/Fall) by @tsutsui : https://speakerdeck.com/tsutsui/kof2016-netbsd-dreamcast-ide-hdd CHAPTER

NINE

OSC2016 KYOTO

• Official page : https://www.ospn.jp/osc2016-kyoto/

• 2016/7/29(Fri), 30(Sat) at Kyoto Research Park (KRP)

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-OSC2016Kyoto.pdf

• Report : http://qml.610t.org/FreeBSD/OSC2016Kyoto.html

• togetter by JNUG : https://togetter.com/li/1005828

• netbsd-advocacy: “NetBSD machines at Open Source Conference 2016 Kyoto” http://mail-index.netbsd.org/netbsd-advocacy/2016/08/01/msg000712.html • LiteBSD at chipKIT Wi-FIRE 4.4BSD PIC32MZ

• RetroBSD at PIC32-RetroBSD 2.11BSD PIC32MX

• Gift box cluster 2 – FreeBSD at Raspberry Pi B FreeBSD LCD loadavg via GPIO

– NetBSD at Banana Pi

– RaspBSD at BeagleBone Black FreeBSD11 LED loadavg, USB Ether+USB Serial via power port – NetBSD at Raspberry Pi 2B LanAP distcc

– NetBSD at Raspberry Pi Zero LED loadavg meter via GPIO

9.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2016kyoto.pdf

• PSG sound at NetBSD/luna : https://speakerdeck.com/tsutsui/osc2016-kyoto-psg-tunes-on-netbsd-luna68k

• RetroBSD & LiteBSD by @610t : http://qml.610t.org/FreeBSD/OSC2016Kyoto_JNUG.html CHAPTER

TEN

OSC2015 KANSAI@KYOTO

• Official Page : https://www.ospn.jp/osc2015-kyoto/

• 2016/8/7(Fri), 8(Sat) at Kyoto Research Park (KRP)

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-OSC2015Kyoto.pdf

• Report : http://qml.610t.org/FreeBSD/OSC2015Kyoto.html

• togetter by JNUG : https://togetter.com/li/857157

• netbsd-advocacy: “OpenBSD and NetBSD machines at Open Source Conference 2015 Kansai@Kyoto” http://mail-index.netbsd.org/netbsd-advocacy/2015/08/10/msg000691.html • Gift box cluster 1: distcc distributed compile cluster

– NetBSD Raspberry Pi B+, 2B, Netwalker • FreeBSD 11 MZTX-PI-EXT LCD via GPIO

10.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2015kyoto.pdf

• Maintainance before 20 years machine at NetBSD : https://speakerdeck.com/tsutsui/osc2015kyoto-netbsd-and-old-machines

CHAPTER

ELEVEN

KOF 2014

• Official page : https://k-of.jp/2014/

• 2014/11/7(Fri), 8(Sat) at ATC (Osaka)

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-KOF2014.pdf

• togetter by JNUG : https://togetter.com/li/742243

• netbsd-advocacy: “NetBSD machines at KANSAI OPEN FORUM 2014” https://mail-index.netbsd.org/netbsd-advocacy/2014/11/08/msg000672.html

11.1 BSD BOF

• NetBSD Guide : http://www.soum.co.jp/~jun/KOF2014.pdf

• LUNA at KOF : https://speakerdeck.com/tsutsui/kof-and-luna-at-netbsd-booth

CHAPTER

TWELVE

OSC2014 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2014-kyoto/

• 2014/8/1(Fri), 2(Sat) at Kyoto Research Park (KRP)

• togetter by JNUG : https://togetter.com/li/700617

• netbsd-advocacy: “NetBSD machines at Open Source Conference 2014 Kyoto” https://mail-index.netbsd.org/netbsd-advocacy/2014/08/09/msg000658.html K*BUG Booth Guide Documentation, Release

12.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2014kyoto.pdf

• NetBSD/luna68k event of the past year : https://speakerdeck.com/tsutsui/osc2014-kansai-at-kyoto-netbsd-luna68k-report

CHAPTER

THIRTEEN

OSC2013 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2013-kyoto/

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-OSC2013Kyoto.pdf • Booth page : http://qml.610t.org/FreeBSD/OSC2013Kyoto.html

• Report : http://qml.610t.org/FreeBSD/OSC2013KyotoReport.html

• togetter by K*BUG : https://togetter.com/li/506220

• togetter by JNUG : https://togetter.com/li/542885 • iCat audio system

• RetroBSD, 2.11BSD on PIC32

• Board, board, board!!

– Squeak Drive a Car with Gainer compatible board (Design Wave Magazine 2008/05:

http://www.cqpub.co.jp/dwm/contents/dwm0126i.htm )

– OpenGL controller with Coldfire board (Interface 2008/09: http://www.cqpub.co.jp/interface/contents/2008/JA/200809.htm )

13.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2013kyoto.pdf

• all about OMRON LUNA, Do LUNA68K Dream of mltermic Twitter Sheep? : http://www.ceres.dti.ne.jp/tsutsui/osc2013kyoto/NetBSD-luna68k_mlterm-fb_Twitter.html

• Board, board, board!! : http://qml.610t.org/FreeBSD/OSC2013Kyoto610.html

CHAPTER

FOURTEEN

EVENT @ NARA NATIONAL COLLEGE OF TECHNOLOGY(NNCT) 13RD

• Official page : http://qml.610t.org/NNCT/event13.html

• 2013/4/6(Sat) at NNCT

• [Hands on] FreeBSD on PC-BSD 9.1

• [Presentation] Let’s enjoy BSD with boards : http://qml.610t.org/FreeBSD/furoku2013.html

CHAPTER

FIFTEEN

KOF 2012

• Official page : https://2012.k-of.jp/ • 2012/11/9(Fri), 10(Sat) at ATC (Osaka)

• More Lights!! Spotlight daemon with OpenBSD/i386 x Gainer

• Squeak+Gainer Drive a Car!!

• Coldfire with 3 axis accelerometer board x OpenBSD/zaurus

• K*BUG POV using Arduino compatible board named Japanino http://otonanokagaku.net/japanino/

15.1 BSD BoF

• NetBSD Guide : http://www.soum.co.jp/~jun/KOF2012.pdf

K*BUG Booth Guide Documentation, Release

• Recent BSD I/O : http://qml.610t.org/FreeBSD/BSD_IO_2012_presen.html

• (What is K*BUG? : http://qml.610t.org/FreeBSD/KBUG_KOF2012_presen.html )

\overline{15.1. BSD BoF 47}

CHAPTER

SIXTEEN

OSC2012 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2012-kyoto/ • 2012/8/3(Fri), 4(Sat) at KRP

• togetter by JNUG : https://togetter.com/li/350035 • RetroBSD, 2.11BSD on PIC32

• USL-5P

• High density zaurus rack.

• More Lights!! Spotlight daemon with NetBSD/hpcmips (WS007SH) x Gainer

16.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2012kyoto.pdf

• Recent NetBSD/luna68k : http://www.ceres.dti.ne.jp/tsutsui/osc2012kyoto/NetBSD-luna68k-updates.html

• Repair power unit of OMRON Luna-II : https://togetter.com/li/354562

CHAPTER

SEVENTEEN

EVENT @ NNCT 12ND

• Official page : http://qml.610t.org/NNCT/event12.html

• 2012/7/29(Sun) at NNCT

• [Hands on] Install PC-BSD http://qml.610t.org/FreeBSD/PCBSD.html • [Booth] Install BSD at various system

• [Presentation] Use BSD at work by BSD-BA : http://www.bsd-ba.org/

CHAPTER

EIGHTEEN

KOF 2011

• Official page : https://k-of.jp/2011/

• 2011/11/11(Fri), 12(Sat) at ATC

• togetter by JNUG : https://togetter.com/li/213724 • Active/Standby Firewall with 2 OpenBSD/landisk at USL-5P : http://www.rururu.org/fswiki/wiki.cgi?page=OpenBSD%2Flandisk%A4%C7Active%2DStandby+Firewall • {Net, Free, Open DragonFly}BSD at NetBSD/xen

18.1 BSD BoF

• NetBSD Guide : http://www.soum.co.jp/~jun/KOF2011.pdf

• Why enigmatic machines? by @tsutsuii : http://www.ceres.dti.ne.jp/tsutsui/kof2011/Why-enigmatic-machines.html

• pkgsrc

CHAPTER

NINETEEN

OSC2011 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2011-kyoto/

• 2011/7/15(Fri), 16(Sat) at KRP

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-OSC2011Kyoto.pdf • OpenBSD/landisk USL-5P

• {Net, Free, Open, DragonFly}BSD at NetBSD/xen

• Visualization NetBSD and FreeBSD build process.

• Squeak+Gainer Drive a Car!! with FreeBSD.

• More Light!! swing daemon based on Arduino.

19.1 JNUG Seminar

• NetBSD Guide : http://www.soum.co.jp/~jun/OSC2011kyoto.pdf

• “NetBSD/m68k will never die!” : http://www.ceres.dti.ne.jp/tsutsui/osc2011kyoto/NetBSD-m68k-will-never-die.html

CHAPTER

TWENTY

OSC2010 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2010-kyoto/

• 2010/7/9(Fri), 10(Sat) at kcg.edu: http://www.kcg.ac.jp/

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-OSC2010Kyoto.pdf

20.1 Booth

• Fastest script for installing FreeBSD.

• Try installation NetBSD/bebox, but fail...

CHAPTER

TWENTYONE

OSC2010 KANSAI@KOBE

• Official page : https://www.ospn.jp/osc2010-kobe/

• 2010/3/13(Sat) at Kobe Industrial Promotion Center

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-OSC2010Kobe.pdf

21.1 JNUG Seminar

• “making of SMC_TT - VME Ethernet using ISA NIC for ATARI TT030” by @tsutsuii : http://www.ceres.dti.ne.jp/tsutsui/osc2010kobe/SMC_TT-OSC2010Kobe.html

• Ethernet card SMC_TT for ATARI TT030 :https://togetter.com/li/76769

CHAPTER

TWENTYTWO

KOF 2009

• Official page : https://k-of.jp/2009/

• 2009/11/6(Fri), 7(Sat) at ATC

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-KOF2009.pdf

• NetBSD/mac68k on LC475Upgrade

CHAPTER

TWENTYTHREE

OSC2009 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2009-kansai/ • 2009/7/10(Fri), 11(Sat) at kcg.edu: http://www.kcg.ac.jp/

23.1 Booth

• LED CPU meter

• BSD daemon statue gadgets

• USL-5P+USB LCD(landisk)

CHAPTER

TWENTYFOUR

NICO-TECH: KYOTO MEETING (NK:M)

• Official page : http://wiki.nicotech.jp/nico_tech/index.php?NTM2%2F%E9%96%A2%E8%A5%BF • 2009/3/21(Sat) at Saiin-Kasuga shrine (Kyoto) http://www.kasuga.or.jp/

• Report : http://qml.610t.org/squeak/mutoh_20090321.html

• Presentation : http://www.nicovideo.jp/watch/sm6562064 (7:00-9:40)

• Squeak-ja meets K*BUG

• On FreeBSD – Squeak+Gainer Drive a Car!!

– Scartch + PicoBoard

– Servo motor daemon’s tail drive by USB audio device

• WorldStethoscope : http://swikis.ddo.jp/WorldStethoscope/6

• Squeak on Pocket Post Pet (Windows CE)

CHAPTER

TWENTYFIVE

KOF 2008

• Official page : https://k-of.jp/2008/

• 2008/11/7(Fri), 8(Sat) at ATC (Osaka)

• K*BUG flyer : http://sacraya.610t.org/kbug/flyer/flyer-KOF2008.pdf

K*BUG Booth Guide Documentation, Release

• OpenBSD/landisk at USL-5P

• Devices: Servo motor deive deamon’s tails using USB audio, Squeak+Gainer

http://www.yengawa.com/squeak_gainer

\overline{73}

CHAPTER

TWENTYSIX

OSC2008 KANSAI@KYOTO

• Official page : https://www.ospn.jp/osc2008-kansai/ • 2008/7/18(Fri), 19(Sat) at kcg.edu: http://www.kcg.ac.jp/

K*BUG Booth Guide Documentation, Release

• BSD for various machines: sgi O2, BeBox(not work), Zaurus, ML115, Fonera, NetBSD/hpcmips, hpcsh, evbmips, alpha (digital Alpha Multia)

• Miku Hatsune paper doll with Six servo moters controlled by USB audio

• Devices: USB one-seg tuner, UHID based themal sensor USB-RH https://strawberry-linux.com/catalog/items?code=52002

CHAPTER

TWENTYSEVEN

KOF 2007

• Official page : https://k-of.jp/2007/ • 2007/11/9(Fri), 10(Sat) at ATC

CHAPTER

TWENTYEIGHT

EVENT @ NNCT 11ST

• Official page : http://qml.610t.org/NNCT/event13.html • 2007/3/31(Sat) at NNCT

• Report : http://qml.610t.org/NNCT/report_event_11th.html K*BUG Booth Guide Documentation, Release

• Introduction NNCT Information Engineering Computer Room.

• FreeBSD ports: RoboCup 2D simulator

84  Chapter 28. Event @ NNCT 11st

K*BUG Booth Guide Documentation, Release

• Create Squeak Dual Boot CDROM : http://qml.610t.org/squeak/squeak_cdrom.html

CHAPTER

TWENTYNINE

KOF 2006

• Official page : https://k-of.jp/2006/ • 2006/11/17(Fri), 18(Sat) at ATC • On FreeBSD

– Croquet: https://en.wikipedia.org/wiki/Croquet_Project

– Squeak + WorldStethoscope

– Moeru Squeak ( http://moesuku.ofg.jp/ ) on ONScripter ( https://onscripter.osdn.jp/onscripter.html )

CHAPTER

THIRTY

EVENT @ NNCT 3RD

• 2002/7/20(Sat) at NNCT

• Report published at FreeBSD Press No12: http://sacraya.610t.org/Press/No12/nnct-event/

30.1 Presentations

• Report: xcast6 online meeting

• Postfix

• One chip microcomputer AVR

• Recovery 48 computers using PXEBOOT in 30 minutes

• Document using SmartDoc

• About DocBook/SGML

CHAPTER

THIRTYONE

EVENT @ NNCT 1ST

• 2001/8/4(Sat) at NNCT

• Report published at FreeBSD Press No7: http://sacraya.610t.org/Press/No7/bugs/

• BSD for various machines.

– NetBSD/hpcmips: NEC MobileGearII 330R, Fujitsu PenCentura130, PocketPostpet – NetBSD/sun3: Sun 3/60

– NetBSD/sun3x: Sun 3/80

– NetBSD/mac68k: Macintosh IIcx

– NetBSD/news68k: Sony NEWS NWS-831,PWS-1550,NWS-1720

• Install hands on

• BSD media in Japan Table of Contents

• What is Kansai * BSD Users Group?

• AsiaBSDCon 2017

• Kansai * BSD Users Group 2017/2/11 meeting

• Open Source Conference (OSC) 2017 Osaka

• KOF 2016

• OSC2016 Kyoto

• OSC2015 Kansai@Kyoto

• KOF 2014

• OSC2014 Kansai@Kyoto

• OSC2013 Kansai@Kyoto

• Event @ Nara National College of Technology(NNCT) 13rd

• KOF 2012

• OSC2012 Kansai@Kyoto

• Event @ NNCT 12nd

• KOF 2011

• OSC2011 Kansai@Kyoto

• OSC2010 Kansai@Kyoto

• OSC2010 Kansai@Kobe

• KOF 2009

• OSC2009 Kansai@Kyoto

• Nico-Tech: Kyoto Meeting (NK:M)

• KOF 2008

• OSC2008 Kansai@Kyoto

• KOF 2007

• Event @ NNCT 11st

• KOF 2006

• Event @ NNCT 3rd

• Event @ NNCT 1st

• Print version PDF is at http://sacraya.610t.org/kbug/book/booth/KBUGBoothGuide-print.pdf

• Valid link PDF is at http://sacraya.610t.org/kbug/book/booth/KBUGBoothGuide.pdf

• HTML version is at http://sacraya.610t.org/kbug/book/booth/singlehtml/

• Source is at https://github.com/610t/kbug/tree/master/book/booth/ K*BUG Booth Guide Documentation, Release 2017/2/25 Photobook edition